CASES99: Logbook Entries

CASES99: Site base Messages, 23 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
202 Thu 10-Feb-2000Pyrgeometer configuration in CASES99basedelany
197 Fri 19-Nov-1999Cosmos configuration for ops4 & Datel problemsbase
196 Fri 19-Nov-1999ops4 (This project won't end!)basemaclean
195 Fri 19-Nov-1999Baud rates of rad data loggersbasemaclean
190 Sat 30-Oct-1999End of Operations basejirak
132 Mon 18-Oct-1999Rebooted cosmos, ragwort, PAM station ingestbasehorst
131 Mon 18-Oct-1999Covars not runningbasehorst
127 Sun 17-Oct-1999CLOVER crashbasesemmer
93 Wed 13-Oct-1999Daily Statusbasehorst
91 Tue 12-Oct-1999writing CDsbasemaclean
90 Tue 12-Oct-1999Hotwire archivebasemaclean
87 Tue 12-Oct-1999How to restart adams & evesbasemaclean
83 Tue 12-Oct-1999Daily Statusbasehorst
76 Mon 11-Oct-1999Daily Statusbasehorst
69 Sun 10-Oct-1999Daily Statusbaseoncley
61 Sat 09-Oct-1999Daily Statusbaseoncley
55 Fri 08-Oct-1999Daily Statusbaseoncley
43 Sun 03-Oct-1999Daily Statusbaseoncley
42 Sat 02-Oct-1999Daily Statusbaseoncley
40 Fri 01-Oct-1999Daily Statusbaseoncley
35 Wed 29-Sep-1999PAM station Configuration Summarybasemilitzer
31 Mon 27-Sep-1999Daily Statusbaseoncley
26 Sat 25-Sep-1999Daily Statusbaseoncley


202: Radiation, Site base, Thu 10-Feb-2000 11:25:28 MST, Pyrgeometer configuration in CASES99
The ten pyrgeoemeters used during CASES99 were set out in two configurations:
 the intercomparison with the pigs mounted side-by side, all up-looking,
 the measurement with the pigs dispersed to their specific location; 
   two up-looking at 50m on the tower boom,
   two down-looking at 50 m on the tower boom,
   two uplooking at 2m, one at station 1, one at station 2,
   four down-looking at 2m, one at station 1, one at station 2, 
                         one at station 3 and one at station 5.
The parameter identifications corresponding to these two configurations are:

           Intercomparison                       Measurement      
            Easterly-most
26416 (Rpile,Tcase,Tdome).out.s02      (Rpile,Tcase,Tdome).out.s2     Bluestem
29260 (Rpile,Tcase,Tdome).in.s02       (Rpile,Tcase,Tdome).in.s2      Bluestem 
31977 (Rpile.Tcase,Tdome).in.s01       (Rpile,Tcase,Tdome).in.s1      Shortgrass
31981 (Rpile,Tcase,Tdome).out.s01      (Rpile,Tcase,Tdome).out.s1     Shortgrass
31975 (Rpile,Tcase,Tdome).out.s05      (Rpile,Tcase,Tdome).out.s5     Weed
31974 (Rpile,Tcase,Tdome).out.a.boom   (Rpile,Tcase,Tdome).out.a.50m  Tower 50m
31980 (Rpile,Tcase,Tdome).in.a.boom    (Rpile,Tcase,Tdome).in.a.50m   Tower 50m
31976 (Rpile,Tcase,Tdome).in.b.boom    (Rpile,Tcase,Tdome).in.b.50m   Tower 50m
31978 (Rpile,Tcase,Tdome).out.b.boom   (Rpile,Tcase,Tdome).out.b.50m  Tower 50m
31979 (Rpile,Tcase,Tdome).out.s03      (Rpile,Tcase,Tdome).out.s3     Drainage
            Westerly-most

The pyrgeometers were in intercomparison configuration:
 1999 September 25, ??:00   to  1999 October 4, ??:00
The pyrgeometers were in measurement configuration :
 1999 October 4, ??:00      to  1999 October 30, ??:00

For more details, including schematics, see:
/net/sssf1/delany/deployments/cases99/lwdiverg/lwdivanal.fm
197: ADAMS, Site base, Fri 19-Nov-1999 16:34:34 MST, Cosmos configuration for ops4 & Datel problems


  Cosmos Matrix card s/n 940211101
   16 channel Ironics serial card s/n 129043
   Datel Dvme 601 A/D card s/n 400190
   signal conditioning card # 3
   NCR power supply s/n 000307
   chassis # 4 (missing insulation on enclosure lid)
   Teca cooler # 4 s/n 8940208


At first we had Datel #440690 in this adam. It would work for
about 1/2 hour and then the A/D data (Tbox.cosmos) would quit.
Saw HSTSTA errors on console once.

Datel #440690 was in daisy in the field.
Swapped in Datel #400190 which was in cosmos in the field.  

The adam failed twice over the weekend in this configuration.
(I mxreset it from home once).
On Monday am (10:00 MST) I plugged the adam's power cable into
a different socket on bottom of the power conditioning box.

At 15:30  MST Monday, replaced coax network cable.  The collision
light on cosmos' transceiver was flickering constantly. Jiggling
the coax connector at the adam end would stop the collisions.












196: BaseSystems, Site base, Fri 19-Nov-1999 16:32:36 MST, ops4 (This project won't end!)

ops4 has been created for the post-intercomparison of the radiometers
on the FLAB roof



195: Radiation, Site base, Fri 19-Nov-1999 16:17:42 MST, Baud rates of rad data loggers
NOTE!

The campbell loggers on the PAM stations transmit at 1200 baud.
They were set this way in the field at CASES99 after the 
intercomparison period.
190: LOG, Site base, Sat 30-Oct-1999 10:48:29 CDT, End of Operations
The end of operations came earlier than we had planned. A significant rain event caused the communications to go down. After talking to Bill Blumen, he said the official end of operations occured 10/30 at 8:00 a.m. We brought up the system in order to archive and shut down the proper way
132: ADAMS, Site base, Mon 18-Oct-1999 10:19:11 CDT, Rebooted cosmos, ragwort, PAM station ingest
Rebooted cosmos and ragwort at 15:18 GMT in order to properly restart
covars.

Also stopped and restartd serial ingest of PAM data on toucan at 
15:20 GMT for the same reason.
131: BaseSystems, Site base, Mon 18-Oct-1999 09:00:07 CDT, Covars not running
The covars quit around 17:25 CDT Oct 17.  Those for daisy and marigold
were restarted following the hangup with Clover, but the others were not 
restarted until 0900 CDT oct 18.
127: BaseSystems, Site base, Sun 17-Oct-1999 19:20:51 CDT, CLOVER crash
While Tom and I were installing the CO2 sensor at 20 meters,
clover suffered a crash. Dave Miller was taking credit for it!
I had to power cycle clover to get it going again. This crash
also shutdown the archiving of blazing data. After a short conversation
with Gordon, marigold and daisy were mxreset to get things going again.

93: Status, Site base, Wed 13-Oct-1999 11:56:06 CDT, Daily Status
ISFF/CASES99 Daily status 13 Oct 1999

Weather: Warm and clear, strong (12m/s) NE winds
Surface: dry
IOPs: IOP6 scheduled for tonight

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      BAD     BAD     up      -       -       -
15m(c)  -       -       -       -       ok      RH      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up
All of the above were down 00:51 - 11:16 today due to a break (rodents?) in
our fiber-optic cable.  We've replaced this cable now so they have more to
chew on.


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok      ok      ok      -       -       ok      ok      ok 	-

General:
- We have rebalanced the AC power load in our trailer and shouldn't suffer
  from further circuit breaker outages.
- We still saw bad signals on stn2 Rpile yesterday, but a cable swap appears
  to have solved it for now.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  We plan to remove this sensor and send it to ATDD soon.
     - The soil heat flux data from stn1 is too large.  I tightened a loose
        connection in the CR10 yesterday, but this didn't appear to help.
RH   - RH values from 15m are low by about 2%, despite replacement of the 
	fan yesterday.	We could replace this sensor with a spare.
[Nap] - We're still waiting for Carmen to connect stn4 to us. 
91: BaseSystems, Site base, Tue 12-Oct-1999 22:03:23 CDT, writing CDs
We write 4 CDs a day:
	stns			~360 Mbytes
	tower			~350 Mbytes
	film 00:00-12:00	~550 Mbytes
	film 12:00-24:00	~550 Mbytes

	total			~1720 Mbytes/day

The archive hard disk is /data/archive/projects/CASES99/raw_data,
an internal disk on cocklebur, which is 8 Gb.

To write a CD, find a cocklebur window, and
put a new CD into the CD-R drive on top of cocklebur

Execute one of the scripts, passing the month number (10=october, duh)
and the day of the month:

	bkcd.stns mon day
	bkcd.tower mon day

first half day of hotfilm data:
	bkcd.film mon day 0 12
last half day:
	bkcd.film mon day 12 24

This script moves files for that day to the "cdbackup" directory,
and then writes them to the CD.  

IF IT SUCCEEDS, IT DELETES FILES FROM THE HARD DISK, so don't
damage the CD. That's your only archive.

The label is then sent to the printer.
Wait for the manual feed light to appear before inserting a label 
sheet. Slide the label sheet into the manual feed slot on
the back of the printer, until it stops, label side toward you,
with an unprinted label in the lower right hand corner.	

After it prints, peel the label of and put it, sticky side up, 
in the CD Stomper. Put the CD, LABEL SIDE DOWN, DATA SIDE UP!!!! 
over the spindle. Anybody who puts the label on the wrong side of
a CD must stay an extra week!!!!

Put the CD in a sleeve, and stick the narrow label on a strip of
perforated paper and slide it into the top slot of the sleeve (use 
the black plastic thingy).

The log of the CD writing is written to $ASTER/projects/CASES99/cdlogs.

Don't write directly on the CD with any pen - reportedly the ink
can leach through and damage the substrate.








	

	







90: BaseSystems, Site base, Tue 12-Oct-1999 21:47:11 CDT, Hotwire archive
A day's hotwire data is 5.0 Gig.

The hotwire files are on
	/data/blaze/aster/projects/CASES99/raw_data/hw

This is an 18 gig disk, so we can hold 3.5 days.  To avoid
running out of space you should archive yesterday's files
each day.

To archive a day's data to disk:

Log onto clover
Insert a new tape into clover's exabyte, then enter:

	blaze2Tape 0 mon day

where mon and day are the month and day to be written to tape.

blaze2Tape takes about 2 hours or so to complete.  It will also delete the
files of (day-1) if they were previously written to tape.

The blaze2Tape log files and tape listings are on:
	$ASTER/projects/CASES99/blazelists

If you experience tape errors,
	insert a "premium" cleaning tape into the drive
	restart the day's archive by deleting the 2 log files for
	that day on $ASTER/projects/CASES99/blazelists,
	and restarting from the beginning of a tape.




87: ADAMS, Site base, Tue 12-Oct-1999 16:49:47 CDT, How to restart adams & eves
Bring up the console window on ASTER so you can watch the adam/eve bootup
messages.

	ping cosmos

If it is alive, you can probably bring it up with mxreset:

	mxreset cosmos

Otherwise you will have to power cycle it.	

In the console window you should see: 

Oct 12 16:52:03 cosmos  System reset!
Oct 12 16:52:29 cosmos  adatsSock SO_SNDBUF len=4096
Oct 12 16:52:29 cosmos  adatsSock SO_SNlast message repeated 1 time
Oct 12 16:52:30 aster cosmos[753]: cosmos Beginning connection on port 1030
Oct 12 16:52:31 cosmos  cosmos booted
Oct 12 16:52:35 aster cosmos[753]: cosmos STREAM socket connected to aster:33008, 1 active connections
etc...etc...

Then do the same for daisy,marigold,ragwort.

If an adam is pingable, but mxreset does not work from ASTER, then try:

	rlogin cosmos
	VxWorks login: aster
	Password: asterize
	-> mxreset

To restart the PAM ingest processes, log into toucan, and do:
	sio_shutdown -s

This should kill 'em all.  If curious, check:
	ps -ef | fgrep stn

If things are OK, then restart them with:
	serial_ingest -z




83: Status, Site base, Tue 12-Oct-1999 10:04:38 CDT, Daily Status
ISFF/CASES99 Daily status 12 Oct 1999

Weather: Warm and clear, S winds at ~5 m/s
Surface: dew this morning, dry by now
IOPs: IOP5 finished this morning (5pm - 8am); none tonight

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp ok
47m(r)  -       -       -       -       -       -       rad ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp ok
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      BAD?    BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      fan?    -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp ok

Data Systems:
ragwort:        up; failed twice yesterday (11:35-13:05 & 14:10-18:25 CDT) 
		due to overheating
marigold:       up
daisy:          up
cosmos:         up
All of the above were down 07:50-08:50 due to a blown circuit breaker.
The blown breaker also caused lost data from the 10m towers 08:15-08:30.


10m TOWERS:
        EVE     Sonic   H2O     Rad     microBar Baro    Prop    T/RH	Soil

1       ok      ok      ok      ok      ok      ok      ok      ok	BAD
2       ok      ok      ok      ok      -       ok      ok      ok	ok
3       ok      ok      ok      ok      -       ok      ok      ok	ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok	-
5       ok      ok      ok      ok      ok      ok      ok      ok 	ok
6       ok      ok      ok      -       -       ok      ok      ok 	-

General:
- An overloaded circuit in our trailer caused power outages through the night.
  We will map out the circuits today to fix this.
- Resolution changed on all 10m kryptons, so we are now able to see smaller
  fluctutations.  We changed to "ops2" in the beginning of IOP5.
- The sonic at 55m was dead through this IOP.  We cycled power this morning
  which revived it.  (We should have done this last night.)
- I don't see noise on the stn6 krypton now.  This may be due to changing
  this station's config not to sample pp.
- We saw a bad signal on stn2 Rpile yesterday, but it appears to be ok now.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 and H2O signals at 20m are noisy, though the average values are now
	reasonable.  Dave Auble found some questionable connections which he
	tightened yesterday.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
WET  - As usual, the CSAT sonic anemometers spike when wet.  
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating and are now connected to
	EVEs at stns 1 and 5.  He hasn't provided a cable to connect to
	stn4 yet.  Our resolution on this signal is about 5 ubars, which may
	be too coarse in some cases (but this is the best we can do).

76: Status, Site base, Mon 11-Oct-1999 11:12:47 CDT, Daily Status
ISFF/CASES99 Daily status 11 Oct 1999

Weather: Warm and clear, SW winds at 5 m/s
Surface: dry
IOPs: ran early this morning (4am - 8am) and another probable for today

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      up      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp up
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      fan?    -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) ok      -       -       ok      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      noisy   -       up      ok      ok      ok 

General:
- All systems ran well last night, though a few sensors are bad.
- All 10m kryptons have coarse resolution, so nighttime data have not been
  well resolved.  We cleaned all kryptons to bring up the signal levels 
  and may change the A/D gain.
- We replaced the sonic at 10m

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 signal at 20m sometimes is noisy and/or goes offscale.  The H2O
	signal from this sensor sometimes is bad also.  We plan to
	bring this sensor down and have Dave Auble look at it in a day or two.
     - Krypton hygrometer at stn6 has some spikes which do not appear in data
       from the other stations.  Cleaning did not change this.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
WET  - As usual, the CSAT sonic anemometers spike when wet.  
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.





69: Status, Site base, Sun 10-Oct-1999 08:45:09 CDT, Daily Status
ISFF/CASES99 Daily status 10 Oct 1999

Weather: Warm, some low-level fog, clear above, light winds
Surface: dew
IOPs: ran last night and another probable for today

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp up
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      ok      -       ok      -
1.5m(c) WET     -       -       ok      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      ok      -       up      ok      ok      ok 

General:
- All systems ran well last night, though a few sensors are bad.
- All 10m kryptons have coarse resolution, so nighttime data have not been
  well resolved.  We'll clean these to bring up the signal levels and may
  change the A/D gain.
- We replaced the T/RH fan at stn3.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 signal at 20m sometimes is noisy and/or goes offscale.  The H2O
	signal from this sensor sometimes is bad also.  We plan to
	bring this sensor down and have Dave Auble look at it in a day or two.
     - Krypton hygrometer at stn6 has some spikes which do not appear in data
       from the other stations.  We'll clean it today.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
WET  - As usual, the CSAT sonic anemometers spike when wet.  In today's case,
	fog has made the 1.5m sensor bad, but others are okay.
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.
61: Status, Site base, Sat 09-Oct-1999 11:21:29 CDT, Daily Status
ISFF/CASES99 Daily status 9 Oct 1999

Weather: Warm, 2/8 fair-weather Cu, light winds
Surface: dry
IOPs: scheduled for tonight and tomorrow

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      ok      -       -       -       -       -
50m(r)  ok      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  ok      ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  ok      -       -       -       -       -       pp up
25m(d)  -       -       -       -       ok      ok      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       ok      ok      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      up      -       ok      -
1.5m(c) ok      -       -       up      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      BAD
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      BAD?    -       up      ok      ok      ok 

General:
- We're finally operational!  Some sensors are not operating properly, but
  most critical sensors should provide good data.  We are ready for this
  IOP.
- daisy now appears to be working
- We replaced the 55m krypton hygrometer and 15m prop which fixed problems
  with these measurements.
- Spikes on CSAT sonic anemometers did disappear when they dried off.
- Spikes on some ATI sonic anemometers may reappear, but probably not in
  light winds.

Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - CO2 signal at 20m sometimes is noisy and/or goes offscale.
     - Krypton hygrometer at stn6 has some spikes which do not appear in data
       from the other stations.
     - The fan on the T/RH sensor at stn3 may not be working.
     - The soil heat flux data from stn1 is too large - probably a bad 
       connection in the CR10.	
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.
55: Status, Site base, Fri 08-Oct-1999 11:47:08 CDT, Daily Status
ISFF/CASES99 Daily status 8 Oct 1999

Weather: Warm, Overcast, light winds, hint of drizzle
Surface: damp
IOPs: possible tomorrow

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  ok      BAD     -       -       -       -       -
50m(r)  WET     -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       ok
45m(r)  -       -       -       -       ok      ok      -
40m(r)  SPIKES  ok      ok      ok      -       -       -
35m(m)  -       -       -       -       ok      ok      -
30m(m)  WET     -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  ok      ok      BAD?    up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  ok      -       -       up      -       -       -
5m(c)   ok      ok      ok      up      -       ok      -
1.5m(c) ok      -       -       up      -       -       pp up

Data Systems:
ragwort:        up
marigold:       up
daisy:          DOWN - NO DATA FROM 15-25m
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       ok      ok      ok      ok      [Nap]   ok      ok      ok
2       ok      ok      ok      ok      up      ok      ok      ok
3       ok      ok      ok      ok      up      ok      ok      ok
4       ok      ok      ok      -       [Nap]   ok      ok      ok
5       ok      ok      ok      ok      [Nap]   ok      ok      ok 
6       ok      ok      BAD?    -       up      ok      ok      ok 


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       been checked.
"ok" - Data appear to be reasonable
BAD  - Data appear to be unreasonable.  We will check and/or replace hygrometers
	at 55m and stn6 today.
[Nap] - Carmen Nappo (NOAA/ATDD) sensors are operating but still not
	cabled into EVEs at stations 1, 4 and 5.
SPIKES - ati.40m has lots of spikes in high winds -- we may fix our spare and
	replace this sensor.
WET   - Campbell sonics often fail with moisture on the transducers.  The higher
	sensors have been in fog all morning, but should resume good operation
	once dry.  

DOWN  - Daisy has recurring problem with crashes -- we are working on this.



43: Status, Site base, Sun 03-Oct-1999 11:50:32 CDT, Daily Status
ISFF/CASES99 Daily status 3 Oct 1999

Weather: Cool, Overcast, lighter winds
Surface: reasonably dry
IOPs: possible tomorrow

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  SPIKES  up      up      up      -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  up      up      up      up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       up      -       -       -
5m(c)   up      up      ready   up      -       up      -
1.5m(c) up      -       -       up      -       -       pp up

Data Systems:
ragwort:        DOWN - NO DATA ABOVE 40m YET
marigold:       up
daisy:          up
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       up      up      up      [IC]    [IC/Nap] up     up      up
2       up      up      up      [IC]    [IC]    up      DOWN    DOWN
3       up      up      up      [IC]    [IC]    up      up      up
4       up      up      up      -       [Nap]   up      up      up
5       up      up      up      [IC]    [Nap]   up      up      up
6       up      up      up      -       up      up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       (yet) been checked.
"ready" - Able to be deployed, but not mounted.  H2O/CO2 sensors at 20m 
	and 40m now mounted and should be connected soon. Licor CO2 at 1.5m
	may be installed today.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
	Since the results appear to shift with time, Jielun and I agree that 
	more cases are necessary.  I cleaned the domes last night which may
	provide some insight.
[Nap] - Carmen Nappo (NOAA/ATDD) is still installing.  These sensors will be
	cabled into EVEs at stations 1, 4 and 5.
SPIKES - ati.40m has lots of spikes in high winds.  I need to call Gary Z. to
	figure out how to increase the transucer gain.
DOWN - T/RH on stn2 not transmitting - a recurring problem which needs to be
       solved.  Also prop died (perhaps after I played with the cable last 
	night).
       Ragwort has recurring problem with crashes.
42: Status, Site base, Sat 02-Oct-1999 09:34:16 CDT, Daily Status
ISFF/CASES99 Daily status 2 Oct 1999

Weather: Cool, Overcast, windy
Surface: reasonably dry - might have been some rain overnight (but not
	before we left at 23:30)
IOPs: none, since rain predicted through the weekend.

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  SPIKES  ready   ready   up      -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  up      ready   ready   up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       up      -       -       -
5m(c)   up      up      ready   up      -       up      -
2.5m(c) DOWN    -       -       up      -       -       pp up

ragwort:        DOWN - we suspect that heating is the problem
marigold:       up
daisy:          up - back to normal (though still some crashes)
cosmos:         up


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       up      up      up      [IC]    [IC]    up      up      up
2       up      up      up      [IC]    [IC]    up      DOWN    DOWN
3       up      up      up      [IC]    [IC]    up      up      up
4       up      up      up      -       [Nap]   up      up      up
5       up      up      up      [IC]    [Nap]   up      up      up
6       up      up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       (yet) been checked.
"ready" - Able to be deployed, but not mounted.  Will try to get the CO2
	sensors installed today.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
	Since the results appear to shift with time, Jielun and I agree that 
	more cases are necessary.  I cleaned the domes last night which may
	provide some insight.
[Nap] - Carmen Nappo (NOAA/ATDD) is still installing.  These sensors will be
	cabled into EVEs at 4 and 5 only.  Their data system will be at 1
	(but we won't get data from it).
SPIKES - ati.40m has lots of spikes in high winds.  I need to call Gary Z. to
	figure out how to increase the transucer gain.
DOWN - T/RH on stn2 not transmitting - a recurring problem which needs to be
       solved.  Also prop died (perhaps after I played with the cable last 
	night).
       csat.1.5m didn't come up after the software upgrade.  We need to figure
	out how to configure it properly.
       Ragwort has recurring problem with crashes.  We're trying to improve
	cooling (either by removing insulation or resetting the thermostat)
	to the box -- also for daisy.

40: Status, Site base, Fri 01-Oct-1999 08:33:21 CDT, Daily Status
ISFF/CASES99 Daily status 1 Oct 1999

Weather: nice.  Clear, not too windy
Surface: reasonably dry
IOPs: none, since rain predicted to start tomorrow through the weekend,
and tonight should be windy

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       pp up
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  up      ready   ready   up      -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       pp up
25m(d)  -       -       -       -       up      up      -
20m(c)  up      [Cook]  [Cook]  up      -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       up      -       -       -
5m(c)   [Mill]  up      ready   up      -       up      -
2.5m(c) [Mill]  -       -       up      -       -       pp up

ragwort:        DOWN - has daisy's parts.  Being worked on today.
marigold:       up
daisy:          DOWN
cosmos:         up. Was down last night by accident


10m TOWERS:
        EVE     Sonic   H2O     Rad     Other   Baro    Prop    T/RH

1       up      up      up      [IC]    [IC]    up      up      up
2       up      up      up      [IC]    [IC]    up      up      DOWN
3       up      up      up      [IC]    [IC]    up      up      up
4       up      up      up      -       [Nap]   up      up      up
5       up      up      up      [IC]    [Nap]   up      up      up
6       up      up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
       (yet) been checked.
"ready" - Able to be deployed, but not mounted.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
	I've started looking at data, but need to do this more.
[Cook] - Awaiting Dave Cook (Argonne) to bring/send one H2O/CO2 sensor.
       The other is in hand, but is missing an AC power cable (but we may just
       plug this into DC anyway)
[Mill] - Dave Miller (Univ. Conn.) is set up except for his 1.5m sonic.  This
       was awaiting a part which arrived yesterday, so should be up today.
[Nap] - Carmen Nappo (NOAA/ATDD) did most installation yesterday - we need
       to be sure that these sensors are cabled into EVE.  Should be up today.

DOWN - T/RH on stn2 not transmitting - a recurring problem which needs to be
       solved.
       Cosmos was left switched off last night.  It is up now.
       Ragwort has recurring problem with crashes.  We're in the middle of
       board swaps to fix.

35: PAM-EVE, Site base, Wed 29-Sep-1999 13:10:10 CDT, PAM station Configuration Summary
The following configurations are available on ASTER
in /home/pam/config and individual station-ready versions
in /home/pam/projects/CASES99/sta_cfg.

On the Rocky they are located in:
c:\pam\cases99\fmt0070
c:\pam\cases99\fmt0071
c:\pam\cases99\fmt0072
c:\pam\cases99\fmt0073
c:\pam\cases99\fmt0074
c:\pam\cases99\fmt0075
c:\pam\cases99\fmt0076

Feature		70	71	72	73	74	75	76
FASTOUT		y	y	y	y	y	y	y
Sync10,despike	y
File metppf	y	y	y	y	y
File /card	y	y	y
File 'cardcopy'				y	y
GPS, timesync	y	y	y	y	y	y	y
stations		1-3,5	4,6	1-3,5	4,6	1-3,5	4,6
logger		y	y		y		y
ATI		y	y	y	y	y	y	y
SBUS		y	y	y	y	y	y	y
power		y	y	y	y	y	y	y
ptb220		y	y	y	y	y	y	y
tsfc		y	y	y	y	y	y	y
< krypton.dat	y	y	y	y	y	y	y


Summary:
	fmt0070	- Used during testing only, not for ops.
		includes sync10 and despike and handy for attempts
		to work on reboot eve code because with 20hz ati,
		a/d and fastout this is good for locking up some 
		stations.

	fmt0071 - Initial implementation includes data storage to
		both ramdisk and /card
	fmt0072 - Same as 71, except the logger is not included.
		use for stations 4,6

	fmt0073 - Primary Config, stations 1,2,3,5
		Basically same as 71 except uses 'cardcopy' feature
		available with eve prom 990926 which does a 1/day 
		card copy instead of writing to it every 5-minutes
	fmt0074 - Primary Config, stations 4,6
		Same as 73 except no logger.

	fmt0075 - Backup contingency if EVE slogs down.
		Minimal config. does not have local data storage.
		IE Do Not Use this if the fastout port to aster and
		archiving is questionable, down, etc.!!
	fmt0076 - same as 75 but no logger

Krypton Declaration Files:
	Two different PAM config files for each krypton hygrometer have
	been created.  One each for sampling with the delta-pressure sensor
	and one for sampling the krypton alone.  The calibrations are
	based on data from Campbell full range, normal conditions.
	The files are called 'kryp13xx.dat' for sensors 1389-1395
	Declaration files are found in directories:
	Without Delta Pressure: for stations 1,2,3:
		On Rocky c:\pam\cases99\krypton\no_dp\
		On aster /home/pam/projects/CASES99/sta_cfg/krypton/no_dp
	With Delta P included for stations 4,5,6
		On Rocky c:\pam\cases99\krypton\with_dp\
		On aster /home/pam/projects/CASES99/sta_cfg/krypton/with_dp


31: Status, Site base, Mon 27-Sep-1999 17:17:41 CDT, Daily Status
ISFF/CASES99 Daily status 27 Sept 1999

Weather: windy, drizzle, cool all day (Severe T-storms last night)
Surface: plants wet, roads muddy, ground still hard
(lots of visitors today: Hatfield, Miller, Ned & Tim, etc.)

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       [Cux]
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  DOWN    [Cook]  [Cook]  [Lund]  -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       [Cux]
25m(d)  -       -       -       -       up      up      -
20m(c)  up      [Cook]  [Cook]  [Lund]  -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       [Lund]  -       -       -
5m(c)   [Mill]  up      ready   [Mill]  -       up      -
2.5m(c) [Mill]  -       -       [Mill]  -       -       [Cux]

ragwort:        up
marigold:       up
daisy:          up
cosmos:         up-one crash


10m TOWERS:
        EVE     Sonic   H2O/CO2 Rad     Other   Baro    Prop    T/RH

1(f)    ?       up      up      [IC]    [IC]    up      up      up
2(g)    ?       up      up      [IC]    [IC]    up      up      up
3(l)    ?       up      spikes  [IC]    [IC]    up      up      up
4(m)    ?       up      up      -       [Nap]   up      up      up
5(p)    ?       up      up      [IC]    [Nap]   up      up      up
6(w)    ?       up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
        been checked.
"ready" - Able to be deployed, but not mounted.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
[Cux] - Awaiting Joan Cuxart (Spain) to bring microbarographs.  We have
        ports (quad-disc probes) ready to be mounted at 30m and 50m.
        (Kurt will do.)
[Lund] - Awaiting Julie Lundquist (CU) to bring hot-wire probes and 
        electronics all of which will need to be mounted.
[Cook] - Awaiting Dave Cook (Argonne) to bring/send H2O/CO2 sensors.
[Mill] - Dave Miller (Univ. Conn.) is now on-site and working on his
	sensor mountings -- a chem shelter has been emptied and moved for him.
[Nap] - Awaiting Carmen Nappo (NOAA/ATDD) to bring sensors.

DOWN - 40m ATI was spiking yesterday, so we removed it.  We swapped stn4's
	ATI with our spare UW to have another ATI to swap back in.  I added
	the hotwire mount to this, but both ATIs are now in the lab.  (It was
	too wet to climb today.)


26: Status, Site base, Sat 25-Sep-1999 13:30:11 CDT, Daily Status
ISFF/CASES99 Daily status 25 Sept 1999

Weather: windy, cloudy in AM, clear midday, overcast in PM
Surface: dry

55m TOWER:
        Sonic   H2O     CO2     HW/F    Prop    T/RH    Other

55m(r)  up      up      -       -       -       -       -
50m(r)  up      -       -       -       -       -       [Cux]
47m(r)  -       -       -       -       -       -       [IC]
45m(r)  -       -       -       -       up      up      -
40m(r)  up      [Cook]  [Cook]  [Lund]  -       -       -
35m(m)  -       -       -       -       up      up      -
30m(m)  up      -       -       -       -       -       [Cux]
25m(d)  -       -       -       -       up      up      -
20m(c)  up      [Cook]  [Cook]  [Lund]  -       -       -
15m(c)  -       -       -       -       up      up      -
10m(c)  up      -       -       [Lund]  -       -       -
5m(c)   up      up      ready   [Mill]  -       up      -
2.5m(c) [Mill]  -       -       [Mill]  -       -       [Cux]

ragwort:        intermittent crashes
marigold:       up
daisy:          up
cosmos:         up-fixed krypton crashes


10m TOWERS:
        EVE     Sonic   H2O/CO2 Rad     Other   Baro    Prop    T/RH

1(f)    ?       up      up      [IC]    [IC]    up      up      up
2(g)    ?       up      up      [IC]    [IC]    up      up      up
3(l)    ?       up      spikes  [IC]    [IC]    up      up      up
4(m)    ?       up      up      -       [Nap]   up      up      up
5(p)    ?       up      up      [IC]    [Nap]   up      up      up
6(w)    ?       up      up      -       [Nap]   up      up      up


Notes:
"up" - Deployed and transmitting data, but quality of operation has not 
        been checked.
"ready" - Able to be deployed, but not mounted.
[IC] - Pyrgeometer intercomparison using all CR10 loggers currently set up
        about 35m S of 55m tower.  Sensors and loggers will be reconfigured
        once we are satisfied that the data from these sensors compare well.
[Cux] - Awaiting Joan Cuxart (Spain) to bring microbarographs.  We have
        ports (quad-disc probes) ready to be mounted at 30m and 50m.
        (Kurt will do.)
[Lund] - Awaiting Julie Lundquist (CU) to bring hot-wire probes and 
        electronics all of which will need to be mounted.
[Cook] - Awaiting Dave Cook (Argonne) to bring/send H2O/CO2 sensors.
[Mill] - Awaiting Dave Miller (Univ. Conn.) to bring sensors.
[Nap] - Awaiting Carmen Nappo (NOAA/ATDD) to bring sensors.